# Tightness for interlacing geometric random walk bridges

### v1

@@ 1.1. Introduction
Part of the interest in ensembles that satisfy the interlacing Gibbs property comes from the fact that they arise in the Schur processes, introduced by Okounkov and Reshetikhin in [OR03]. In particular, in Section 6 we apply Theorem 1.1 to a sequence of spiked Schur processes, that were recently considered in [Dim24], and show that the latter converge uniformly over compact sets to the Airy wanderer line ensembles – see Proposition 6.2. We mention that the Airy wanderer line ensembles were constructed for general parameters in [Dim24] as weak limits of such ensembles for special parameters from [AFVM10, CH14]. By combining [Dim24, Theorem 1.8] and Proposition 6.2 one obtains an alternative construction of these ensembles, directly as weak limits of Schur processes, which is independent of the results in [AFVM10, CH14]. In addition, the more general Theorem 5.1 in the main text allows one to deal with line ensembles defined on a half-infinite or even a finite interval, which makes it applicable to a large class of models including the half-space and free boundary Schur processes from [BBCS18, BBNV18], and the periodic Schur processes from [BB19, Bor07]. In the upcoming work [DY24] we use the results of this paper to construct half-space analogues of the Airy line ensemble as weak limits of the half-space Schur processes.
@@ 1.2. Main result
Theorem 1.1, and more generally Theorem 5.1, is similar to [CH14, Theorem 3.8], [DFF+21, Theorem 1.1], and [Ser23, Theorem 1.1]. In [DFF+21] the authors dealt with avoiding Bernoulli random walk bridges, and [Ser23] extended the argument to a class of avoiding random walkers with general jump distributions, while [CH14] worked directly with avoiding Brownian line ensembles. One of the differences between our setup and [DFF+21, Ser23] is that our walkers are interlacing rather than avoiding – this is quite mild and does not affect the problem substantially. A more serious difference is that in [DFF+21, Ser23] the authors only worked on $\mathbb{R}$, assumed that the top curves $\mathcal{L}_{1}^{N}$ have one-point tight marginals, which are globally parabolic. This global parabolicity is what allowed them to conclude that all the other curves also need to have tight one-point marginals. In Theorem 5.1 we want to deal with possibly finite intervals $\Lambda$, for which one needs more than just information about the top curve. Specifically, it is possible for all curves except the first ones in our sequence of ensembles to dip to $\infty$- ∞, and the top curves to converge to a Brownian bridge on $\Lambda$. Consequently, one-point tightness of the top curves does not prevent other curves from escaping to $\infty$- ∞. This is why we require one-point tightness of all the curves and not just the top one. In this regard our assumptions are closer to those in [CH14].
@@ 1.3. Outline
In Section 2 we introduce the general class of models we study, called geometric line ensembles, and their local interlacing structure, called the interlacing Gibbs property. The section contains most of the core notation for the rest of the paper, as well as the monotone coupling, see Lemma 2.10; the comparison to Brownian bridges, see Proposition 2.11; and the statement that under diffuse scaling interlacing geometric random walk bridges converge to avoiding Brownian bridges, see Lemma 2.14. The proofs of Lemmas 2.10 and 2.14 are given in Appendix A. Section 3 provides several technical estimates for a single geometric random walk bridge, which is either free, or conditioned to stay above a well-behaved lower boundary. The proofs in Section 3 rely mostly on the monotone coupling in Lemma 2.10 and the bridge comparison in Proposition 2.11. Section 4 contains the critical Lemma 4.1, which ensures curve separation, and Lemma 4.3, which provides modulus of continuity estimates under the assumption of curve separation. Section 5 contains the most general version of our results, see Theorem 5.1. In the same section we deduce Theorem 1.1 from Theorem 5.1 and provide a proof of the latter. In Section 6 we apply Theorem 1.1 to sequences of line ensembles that arises in spiked Schur processes, and show that the latter converge to the Airy wanderer line ensembles, see Proposition 6.2.
@@ 2.1. Geometric line ensembles
Given two integers $a\leq b$, we let $\llbracket a,b\rrbracket$ denote the set $\{a,a+1,\dots,b\}$. We also set $\llbracket a,b\rrbracket=\emptyset$ when $a>b$, $\llbracket a,\infty\rrbracket=\{a,a+1,a+2,\dots\}$, $\llbracket-\infty,b\rrbracket=\{b,b-1,b-2,\dots\}$ and $\llbracket-\infty,\infty\rrbracket=\mathbb{Z}$. Given an interval $\Lambda\subset\mathbb{R}$, we endow it with the subspace topology of the usual topology on $\mathbb{R}$. We let $(C(\Lambda),\mathcal{C})$ denote the space of continuous functions $f:\Lambda\rightarrow\mathbb{R}$ with the topology of uniform convergence over compacts, see [Mun03, Chapter 7, Section 46], and Borel $\sigma$-algebra $\mathcal{C}$. Given a set $\Sigma\subset\mathbb{Z}$, we endow it with the discrete topology and denote by $\Sigma\times\Lambda$ the set of all pairs $(i,x)$ with $i\in\Sigma$ and $x\in\Lambda$ with the product topology. We also denote by $\left(C(\Sigma\times\Lambda),\mathcal{C}_{\Sigma}\right)$ the space of real-valued continuous functions on $\Sigma\times\Lambda$ with the topology of uniform convergence over compact sets and Borel $\sigma$-algebra $\mathcal{C}_{\Sigma}$. We typically take $\Sigma=\llbracket 1,N\rrbracket$ with $N\in\mathbb{N}\cup\{\infty\}$.
@@
and for any finite $K=\llbracket k_{1},k_{2}\rrbracket\subset\llbracket 1,N-1\rrbracket$ and $a,b\in\llbracket T_{0},T_{1}\rrbracket$ with $a<b$ the following holds. Suppose that $f,g$ are two increasing paths drawn in $\{(r,z)\in\mathbb{Z}^{2}:a\leq r\leq b\}$ and $\vec{x},\vec{y}\in\mathfrak{W}_{k}$ with $k=k_{2}-k_{1}+1$ altogether satisfy that $\mathbb{P}(A)>0$ where $A$ denotes the event
@@
where if $k_{1}=1$ we adopt the convention $f=\infty=L_{0}$. Then, writing $k=k_{2}-k_{1}+1$, we have for any $\{B_{i}\in\Omega(a,b,x_{i},y_{i})\}_{i=1}^{k}$ that
@@
Observe that in Definition 2.6 the index $k_{2}$ is assumed to be less than or equal to $N-1$, so that if $N<\infty$ the $N$-th path is special and is not conditionally uniform. If $N=1$, then the conditions in Definition 2.6 become void, i.e., any geometric line ensemble with one line satisfies the geometric Gibbs property. We mention that since $\mathfrak{L}$ is interlacing we automatically have for each $j\in\llbracket T_{0},T_{1}\rrbracket$ that $L_{1}(j)\geq L_{2}(j)\geq\cdots$. In addition, the well-posedness of $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ in (2.3) is a consequence of our assumption that $\mathbb{P}(A)>0$.

### v2

@@ 1.1. Introduction
Part of the interest in ensembles that satisfy the interlacing Gibbs property comes from the fact that they arise in the Schur processes, introduced by Okounkov and Reshetikhin in [OR03]. In particular, in Section 6 we apply Theorem 1.1 to a sequence of spiked Schur processes, that were recently considered in [Dim24], and show that the latter converge uniformly over compact sets to the Airy wanderer line ensembles – see Proposition 6.2. We mention that the Airy wanderer line ensembles were constructed for general parameters in [Dim24] as weak limits of such ensembles for special parameters from [AFVM10, CH14]. By combining [Dim24, Theorem 1.8] and Proposition 6.2 one obtains an alternative construction of these ensembles, directly as weak limits of Schur processes, which is independent of the results in [AFVM10, CH14]. In addition, the more general Theorem 5.1 in the main text allows one to deal with line ensembles defined on a half-infinite or even a finite interval, which makes it applicable to a large class of models including the half-space and free boundary Schur processes from [BBCS18, BBNV18], and the periodic Schur processes from [BB19, Bor07]. For example, the results of the present paper have already found applications to the analysis of half-space Schur processes away from the boundary in [Zho25] and in constructing half-space analogues of the Airy line ensemble in [DY25].
@@ 1.2. Main result
Theorem 1.1, and more generally Theorem 5.1, is similar to [CH14, Theorem 3.8], [DFF+21, Theorem 1.1], and [Ser23, Theorem 1.1]. In [DFF+21] the authors dealt with avoiding Bernoulli random walk bridges, and [Ser23] extended the argument to a class of avoiding random walkers with general jump distributions, while [CH14] worked directly with avoiding Brownian line ensembles. One of the differences between our setup and [DFF+21, Ser23] is that our walkers are interlacing rather than avoiding – this is quite mild and does not affect the problem substantially. A more serious difference is that in [DFF+21, Ser23] the authors only worked on $\mathbb{R}$, assumed that the top curves $\mathcal{L}_{1}^{N}$ have one-point tight marginals, which are globally parabolic. This global parabolicity is what allowed them to conclude that all the other curves also need to have tight one-point marginals. In Theorem 5.1 we want to deal with possibly finite intervals $\Lambda$, for which one needs more than just information about the top curve. Specifically, it is possible for all curves except the first ones in our sequence of ensembles to dip to -$\infty$, and the top curves to converge to a Brownian bridge on $\Lambda$. Consequently, one-point tightness of the top curves does not prevent other curves from escaping to -$\infty$. This is why we require one-point tightness of all the curves and not just the top one. In this regard our assumptions are closer to those in [CH14].
@@ 1.3. Outline
In Section 2 we introduce the general class of models we study, called geometric line ensembles, and their local interlacing structure, called the interlacing Gibbs property. The section contains most of the core notation for the rest of the paper, as well as the monotone coupling, see Lemma 2.12; the comparison to Brownian bridges, see Proposition 2.13; and the statement that under diffuse scaling interlacing geometric random walk bridges converge to avoiding Brownian bridges, see Lemma 2.16. The proofs of Lemmas 2.12 and 2.16 are given in Appendix A. Section 3 provides several technical estimates for a single geometric random walk bridge, which is either free, or conditioned to stay above a well-behaved lower boundary. The proofs in Section 3 rely mostly on the monotone coupling in Lemma 2.12 and the bridge comparison in Proposition 2.13. Section 4 contains the critical Lemma 4.1, which ensures curve separation, and Lemma 4.3, which provides modulus of continuity estimates under the assumption of curve separation. Section 5 contains the most general version of our results, see Theorem 5.1. In the same section we deduce Theorem 1.1 from Theorem 5.1 and provide a proof of the latter. In Section 6 we apply Theorem 1.1 to sequences of line ensembles that arises in spiked Schur processes, and show that the latter converge to the Airy wanderer line ensembles, see Proposition 6.2.
@@ 2.1. Geometric line ensembles
[…]
@@
and for any finite $K=\llbracket k_{1},k_{2}\rrbracket\subseteq\llbracket 1,N-1\rrbracket$ and $a,b\in\llbracket T_{0},T_{1}\rrbracket$ with $a<b$ the following holds. Suppose that $f,g$ are two increasing paths drawn in $\{(r,z)\in\mathbb{Z}^{2}:a\leq r\leq b\}$ and $\vec{x},\vec{y}\in\mathfrak{W}_{k}$ with $k=k_{2}-k_{1}+1$ altogether satisfy that $\mathbb{P}(A)>0$ where $A$ denotes the event
@@
where if $k_{1}=1$ we adopt the convention $f=\infty=L_{0}$. Then, we have for any $\{B_{i}\in\Omega(a,b,x_{i},y_{i})\}_{i=1}^{k}$ that
@@
Observe that in Definition 2.6 the index $k_{2}$ is assumed to be less than or equal to N1N-1, so that if $N<\infty$ the $N$-th path is special and is not conditionally uniform. If $N=1$, then the conditions in Definition 2.6 become void, i.e., any geometric line ensemble with one line satisfies the geometric Gibbs property. We mention that since $\mathfrak{L}$ is interlacing we automatically have for each $j\in\llbracket T_{0},T_{1}\rrbracket$ that $L_{1}(j)\geq L_{2}(j)\geq\cdots$. In addition, the well-posedness of $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ in (2.3) is a consequence of our assumption that $\mathbb{P}(A)>0$.
@@
We end this section with the following two results. The first, Lemma 2.10, shows that line ensembles with law $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},\infty,g}$ as in Definition 2.5 satisfy the interlacing Gibbs property. The second, Lemma 2.11, provides a stronger property than (2.3) enjoyed by line ensembles that satisfy the interlacing Gibbs property. This condition is automatically satisfied; however, it is a bit harder to check in practice, which is why we formulated the interlacing Gibbs property using (2.3). Both lemmas are proved in Appendix A.3.
Lemma 2.10.
Let $\vec{x},\vec{y}\in\mathfrak{W}_{k}$, $T_{0},T_{1}\in\mathbb{Z}$ with $T_{0}<T_{1}$, and $g:\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$ be an increasing path. Suppose that $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x},\vec{y},\infty,g)\neq\emptyset$ and $\mathfrak{Q}$ has law $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},\infty,g}$ as in Definition 2.6. Define the $\llbracket 1,k+1\rrbracket$-indexed geometric line ensemble $\mathfrak{L}:\llbracket 1,k+1\rrbracket\times\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$ through
$\mathfrak{L}(i,j)=\mathfrak{Q}(i,j)\mbox{ for }i\in\llbracket 1,k\rrbracket,j\in\llbracket T_{0},T_{1}\rrbracket,\mbox{ and }\mathfrak{L}(k+1,j)=g(j)\mbox{ for }j\in\llbracket T_{0},T_{1}\rrbracket.$
Then, $\mathfrak{L}$ satisfies the interlacing Gibbs property from Definition 2.6 with $N=k+1$.
Lemma 2.11.
Assume the same notation as in Definition 2.6. Then, for any $\{B_{i}\in\Omega(a,b,x_{i},y_{i})\}_{i=1}^{k}$
(2.4) ${\bf 1}_{A}\cdot\mathbb{P}\left(L_{i+k_{1}-1}\llbracket a,b\rrbracket=B_{i}\mbox{ for $i\in\llbracket 1,k\rrbracket$}\,|\mathcal{F}_{\mathrm{ext}}\right)={\bf 1}_{A}\cdot\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{a,b,\vec{x},\vec{y},f,g}\left(\cap_{i=1}^{k}\{Q_{i}=B_{i}\}\right),$
where $\mathcal{F}_{\mathrm{ext}}=\sigma(L_{i}(j):(i,j)\in\Sigma\times\llbracket T_{0},T_{1}\rrbracket\setminus\llbracket k_{1},k_{2}\rrbracket\times\llbracket a+1,b-1\rrbracket)$.